Such units refer to the diag of the screen, A 25 in t.v set also has a width of a 20 inches what is the height
Answer by nerdybill(5404)

Such units refer to the diag of the screen, A 25 in t.v set also has a width of a 20 inches what is the height
.
The "width", "height", and "diagonal" edge of the TV forms a "right triangle". Because of this, we can apply Pythagorean theorem:
width^2 + height^2 = diagonal^2
.
Plugging in our given numbers:
20^2 + height^2 = 25^2
height^2 = 25^2 - 20^2
height^2 = 625 - 400
height^2 = 225
height = sqrt(225)
height = 15 inches